Two identical mechanical waves travel different distance to a place where they interfere. One ware travels the distance of 20m, and second wave the distance of 60m. Wave length is the same 30 m. Calculate the phase offset, an inference ( constructive, or destructive).
1
Expert's answer
2020-09-23T08:12:36-0400

path difference =30m

"\Delta" = constructive inference ="n\lambda"

=destructive inference ="\dfrac{2n+1}{2}\lambda"

hence "\Delta" =30m

= "\dfrac{3}{2}*20"

="\dfrac{2(1)+1}{2}\lambda" ="\dfrac{2n+1}{2}\lambda"

hence there will be destructive wave inference